OpenStudy (anonymous):

An astronaut weighs 174 pounds on Earth and 29 pounds on the Moon. If his daughter weighs 108 pounds on Earth, what is the daughter's weight on the moon, in pounds?

OpenStudy (austinl):

Set up a proportion. \(\displaystyle \frac{174}{29}=\frac{108}{x}\) Solve for x.